Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
lumaca, rocchi, valentini - 19:03 Saturday 23 October 2021 (53622) Print this report
ISC shift report: carm null recovery and TCS tuning

The aim of the shift was dedicated to checking the stability of the CARM null status after yesterday TCS tuning operation and in supporting further tuning activities (see dedicated entry).

At ~11.30 UTC we lost connection with two DAQboxes of SDB2 and  Pacaud intervened to restore it: 53619


At the beginning of the shift the ITF was locked and brought to CARM offset zero three times (with mild difficulties in achieving stable locks of the CITF) to check the stability of the current TCS configuration. One lock lasted only a few seconds, while the other two lasted ~8 minutes, and showed a quick degradation of the sidebands power that ultimate led to unlocks. Figure 1 shows a comparison between the first (continuous line) and the third  (dotted line) carm offset zero attempts.

The TCS configuration (Both DAS and CH powers) was therefore changed (see dedicated entry) at ~12.57 UTC while the CITF was locked. (See dedicated entry for details). The change resulted in a noticeable loss of sideband power (see figure 2) in the CITF configuration. Attempts of reducing the CARM offset after this tuning immediately failed (before the CARM to MC handoff), most likely due to the gain loss of the CITF DOFs, in particular when locked to the 3f. 

The CH was values were therfore changed at 15.20 UTC  (see figure 3) and allowed to partially recover the sidebands values. Since the 112 MHz was not recovered completely, the MICH and SRCL gain have been increased in DRMI_lock.ini from 2.5 to 3.5 and from 0.4 to 0.6 respectively in order to restore their UGF.

A last attempt to go to CARM offset zero (to evaluate the new DAS tuning) was performed at ~16.20 UTC but the ITF unlocked just before reaching the carm offset zero state.

The ITF was left with the arms locked on the IR beam.


Images attached to this report
Search Help